# End User Authoring of Personalized Content Classifiers: Comparing Example Labeling, Rule Writing, and LLM Prompting

### v1

@@ 3. Methods
We conducted a within-subjects experiment with 37 non-programmer participants to comparatively evaluate these three personal moderation systems for everyday users. Participants used each of the systems—in a randomized order—to build a classifier within a fixed amount of time, resulting in three experimental conditions. Throughout the duration of each condition, we collected the performance of participants’ classifiers every half-minute in order to understand how well each system enables rapid initialization and iterative improvement over time. After each condition, we collected participants’ self-reported usability ratings of the system in that condition to understand their experiences and challenges that they encountered. For 13 participants, we additionally conducted a semi-structured interview after the experiment was over to understand their challenges in creating and iterating personal classifiers. We further describe our system implementations and experiment design below.
@@ 3.2.1. Recruitment and Participants
We recruited 37 non-programmer participants by advertising a call for participation on mailing lists of non-Computer Science departments at two major U.S.-based universities. We only selected participants who self-reported having little knowledge of programming and algorithms. There were 27 females, 9 males, and one participant who preferred not to disclose their gender. Most participants were pursuing their bachelor’s degrees except four pursuing more advanced degrees. Regarding political stance, 16 participants identified as liberal, 12 as moderate, and the remaining 9 were evenly distributed among “very liberal,” “conservative,” and “prefer not to disclose” groups. In terms of generative AI usage, 14 participants used it at least every few weeks, 14 participants every few months, and 9 participants rarely.
@@ 3.3.1. Study Design and Procedure
We employed a within-subjects design with the three experiment systems described above. The final experiment protocol was designed iteratively through five pilot experiments to ensure its effectiveness. This study was reviewed by our university IRB and deemed exempt.
@@ 3.3.2. Evaluation Measures
3.4. Data Analysis
3.4.1. Quantitative Modeling
Quantitative results were analyzed using a linear mixed-effects (LME) model, where experiment systems were treated as a fixed effect and participants as a random effect. The dependent variables in our model included classification performance, creation speed, and various subjective measures of user experience. We did not include the order in which participants used the three systems as another fixed effect, as no significant differences were observed for this variable. For each dependent variable, we calculated pairwise differences between the three experiment systems. Finally, we conducted a sanity check to confirm that participants exhibited diverse moderation preferences. We found that half of the ground truth comments had, at most, a 75% majority consensus.
3.4.2. Qualitative Coding
4. Results
@@ 4.1. What strategies enabled participants to create high-performing classifiers quickly?
Figure 6 compares classification performances across the three systems over 15 minutes, with each line’s endpoint representing the final performance. Table 1 presents the pairwise differences in the final performances among the three systems. We observe no significant differences in accuracy across systems. In terms of precision, the Rule System showed significantly higher precision than the Label System (Est.Diff = 0.097, $p$ 0.05) but did not differ significantly from the Prompt System. However, the Prompt System demonstrated significantly higher recall compared to both the Label System (Est.Diff = 0.205, $p$ 0.001) and the Rule System (Est.Diff = 0.232, $p$ 0.001). Consequently, the Prompt System also achieved a significantly higher $F_{1}$ score than the other two systems (Est.Diff compared to the Rule System = 0.105, $p$ 0.01; Est.Diff compared to the Label System = 0.141, $p$ 0.001). We focus on $F_{1}$ as a performance metric of interest over accuracy due to the unbalanced nature of participants’ labels. Finally, when examining how the final performance of each system varied across participants, we found that the Prompt System showed the least variance, regardless of the performance metric used. This suggests that the Prompt System is less susceptible to individual variances than the other systems.
@@
Table 2 and Table 3 present the pairwise differences in classifier performance at 5 and 10 minutes respectively. Consistent with our earlier findings, the Prompt System demonstrated significantly higher recall and $F_{1}$ scores than the other two systems. Notably, its recall at 5 minutes had already surpassed the final recall of the other two systems. Meanwhile, the Prompt System’s recall exhibited a distinct pattern: a rapid increase in the first five minutes, followed by a steady climb in the subsequent five minutes, and a plateau in the remaining time. In contrast, the Rule System showed a consistent increase in recall throughout the entire 15-minute period, whereas the Label System showed a less noticeably upward trend in recall over time. On average, participants reached 95% of their peak recall by 7 minutes, and that of their peak $F_{1}$ score by 5 minutes. We also observed that participants made fewer changes to their prompts as time went on, while they continued actively labeling examples or curating keywords for their rules in the other two systems. These findings suggest that writing prompts can facilitate rapid initialization but may be less effective in supporting iterative improvements.
@@
The precision of the Prompt System and the Rule System remained comparable throughout the experiment. While the Prompt System showed significantly higher precision than the Label System at 5 minutes (Est.Diff = 0.131, $p$ 0.05), the difference became non-significant at 10 minutes (Est.Diff = 0.082, $p$ = 0.090). In contrast, while initially comparable at 5 minutes (Est.Diff = 0.100, $p$ = 0.06), the Rule System had a significantly higher precision than the Label System at 10 minutes (Est.Diff = 0.099, $p$ 0.05) and at the end (Est.Diff = 0.097, $p$ 0.05).
While both the Prompt System and the Rule System showed notable improvements in recall by the end of the experiment, their precision either plateaued or even decreased over time. For instance, the Rule System experienced a gradual decrease in precision during the first 5 minutes, followed by a steady state for the remaining 10 minutes. The Prompt System, despite a slight increase in precision in the initial 2 minutes, subsequently experienced a minor decline before reaching a plateau. These observations align with participant behaviors: participants tended to add more categories of unwanted content rather than increase the specificity of existing rules or prompts during the task because the latter was often considered more cognitively demanding. Although the Label System demonstrated a rapid increase in precision during the first five minutes, it began with such low precision that its final precision was still below that of the other two systems.
@@ 4.2. What strategies did participants find easiest for communicating their preferences?
Unlike professional moderators, social media users often lack the time or opportunity to articulate their moderation preferences clearly (Jhaver et al., 2023). Instead, their moderation decisions are often a result of their intuitive feelings about the content they encounter. P4 pointed out that “sometimes it might be the entire sentence that I don’t want to see, or maybe it’s just about a word…It could also be the feeling of the comment. You don’t feel right about it.” Some participants also lacked “a holistic view about what were the alarming comments” (P2), since new content continuously flooded their social media feeds. Labeling examples thus provided them with a natural way to express their intuitive preferences to the algorithm, whereas writing prompts or authoring rules forced them to distill their intuitive preferences into high-level patterns. Participants tended to trust algorithms to infer their preferences from their labels more than they trusted themselves to accurately convey their preferences, especially in time- and attention-constrained settings of personal moderation. As P2 explained, “I would actually trust the algorithm finding trends more [than] me looking at everything and trying to build my own…especially with my lack of experience and lack of time.” This trust made participants more willing to offload the task of summarizing high-level criteria to the algorithm, despite this requiring them to label many examples.
@@ 4.3. What strategies were easiest for iterating on classifiers after initialization?
Participants had a clearer understanding of why the Rule System made incorrect predictions than the other two systems. Such transparency helped participants improve their classifiers in some cases, like when they forgot to include spelling variants or similar phrases to those that they explicitly included in their rules. But in other cases, transparency did not lead to iterative improvements because rules could not accommodate contextual variances, as documented in prior research (Song et al., 2023; Reddit, 2023). Our survey and interviews suggest that participants had a slightly better understanding of why their classifiers made mistakes for the Prompt System than for the Label System (Figure 7). Participants found that they could at least review and reflect on problematic prompts in the Prompt System, whereas they could not pinpoint specific labels that led to classification mistakes in the Label System. P2 noted this advantage: “The prompt system was more transparent in the sense that it highlights specific guidelines you have set up. You could then change them and see what they do.
For the Label System, participants had only a vague understanding of why their classifiers made incorrect predictions, reflecting previous research on how end users develop folk theories to interpret content curation algorithms (Eslami et al., 2016; Mayworm et al., 2024; DeVito et al., 2017). Moreover, it offered few opportunities for targeted iteration, which were particularly useful as participants found it easier to point out what the classifier should learn from its mistake. This highlights the need for more flexible approaches at different stages of classifier creation, as recent studies advocated (Piorkowski et al., 2023; Daly et al., 2021) In the absence of such features in the current system, participants could only label more examples similar to these mistakes in the hope of providing targeted feedback.
@@
While participants could quickly build classifiers with decent performance using the Prompt System, they found it challenging to incorporate more nuances when iterating on their prompts. Participants rated it as comparably difficult to correct with the Rule System (Figure 7). Many participants described a disconnect in how humans and LLMs perceive the same prompts. They observed that “descriptions could be very subjective” (P4) and asked, “What has [the LLM] previously been trained on? For words they labeled as extreme, I might not think as extreme” (P13). Participants also suggested that LLMs might not fully understand participants’ experiences and personalities from their prompts and might therefore fail to execute their preferences faithfully. Consequently, many noticed that LLMs seemed to interpret their preferences broadly, causing LLMs to have trouble distinguishing different degrees of a concept. For instance, P13 noted that “[LLMs] have the problem of differentiating between just minor attacks on someone’s character [such as calling someone a fool] and ones that are strong and can be perceived as harm.” Similarly, P8 struggled to teach LLMs to “differentiate between threats and just using the word ‘shoot’ or ‘kill.’” Articulating the fine-grained differences between concepts proved too difficult for many participants. While they could easily teach LLMs what was clearly good or bad, aligning LLMs with their moderation decisions for borderline cases remained a challenge.
@@ 4.4. Which strategies did participants prefer overall?
We asked participants to rank the three systems regarding their preferences for using them in real-life content moderation and to explain their rationale for these rankings. Figure 8 illustrates the distribution of participants’ rankings. Surprisingly, despite growing enthusiasm for using LLMs to moderate content, a comparable number of participants still preferred labeling examples or authoring rules to create custom classifiers for content moderation. While the Prompt System was rated first most often, it was also rated third most often, suggesting that it was a polarizing choice. The Wilcoxon Signed-Rank test showed no significant differences between any of the systems.
@@
  • Whether users want explicit articulation of their preferences. Individual preferences for content curation are nuanced and subject to change. During the experiment, many participants tended to modify their criteria after reviewing more examples. As a result, if participants had explicitly stated their preferences in rules or prompts, they would have to continually reflect on their preferences and update their classifiers accordingly. In comparison, the Label System’s algorithms had the potential to learn these subtle changes in preferences from a continuous flux of labels. As P12 explained, “Preferences change over time, and we may not explicitly realize that. So you may accidentally remove some things if your preferences change but you have already explicitly set them in stone. ” However, some participants appreciated the chance to actively reflect on and articulate their preferences through high-level rules or prompts, despite the additional effort required.
@@
  • Whether users can tolerate reviewing many toxic examples. In our study, participants reviewed fewer comments to develop an effective classifier using the Prompt System compared to the other two systems. On average, participants reviewed 133 and 127 examples for the Label System and the Rule System respectively, whereas only 53 examples were needed for the Prompt System. The need to review numerous toxic examples could be distressing, particularly for those creating classifiers to block such content. Whether people are willing to undertake such an emotional toll can depend on their sense of responsibility in specific moderation scenarios. As P8 stated, “If I were moderating for my own sake and not for a community, I would probably just focus more on the ease of creating a filter so that I don’t have to subject myself to these potentially offensive comments.
@@
  • Whether users prioritize precision over recall. Participants demonstrated varied approaches to balancing precision and recall. Some favored precision because they were more concerned about the risk of removing benign comments than approving unwanted content. P12 described their rationale as follows: “My preference is to limit the number [of comments] that accidentally gets removed. I don’t want to fall into any echo chamber. Even if there are really crazy people saying really crazy things, they make me frustrated, but I still should know about it.” Other participants favored recall, showing a stronger inclination to remove toxic comments, even if it meant mistakenly removing benign content. P13 expressed this preference: “Personally, too specific is worse than the too broad. If the filter keeps certain content I really don’t want to see, that’s far worse than if I didn’t get an extra post. Because of how much content there already is, I’ll just get others on my feed anyway.” For these participants, the high recall of the Prompt System is even advantageous during the creation stage. Since LLMs could collect the majority of potentially unwanted examples given initial criteria, participants could focus on reviewing these examples to increase the precision of their prompts.
@@
  • To what extent users value transparency and controllability. An important distinction between the Rule System and the other two systems lies in transparency. As P10 described, “I like rules more because I know what’s going into them. But with the prompts, you don’t know what’s going on. While the prompting is better at catching things, I just like the transparency of rules more.” The Rule System helps users not only understand why exactly a comment is removed; it also enables moderators and users to have clear expectations about which comments will be removed. For users who moderate for a community, this sense of controllability could enable them to be more accountable for their moderation decisions. For example, P9 shared, “I think there’s like a fine line between filtering or censoring. I think it’s important to know why certain comments or text was filtered.”
@@ 5. Discussion
In this work, we compared three prominent strategies for creating custom classifiers, focusing on their support for rapid initialization and easy iteration. Our experiments revealed that writing prompts generally enabled participants to create custom classifiers with the highest performance most quickly. However, this approach also had shortcomings, such as challenges in further iteration and lack of transparency. In this section, we discuss how incorporating labeling examples and authoring rules could help mitigate these problems, and how the unique user needs of diverse content curation applications demand different kinds of hybrid systems.
@@ 5.1. Hybrid Approaches to Personalized Classifier Creation
Existing tools often fail to accommodate the diversity and fluidity of user needs in content curation. While writing prompts generally proves to be an effective way to convey user preferences to classifiers, our experiment indicates the potential to incorporate labeling examples and authoring rules into future content curation systems, thereby providing end users with more flexible strategies to create and iterate on their classifiers.

### v2

@@ 3. Methods
We conducted a within-subjects experiment with 37 participants with no programming experience to comparatively evaluate these three personal moderation systems for everyday users. Participants used each of the systems—in a randomized order—to build a classifier within a fixed amount of time, resulting in three experimental conditions. Throughout the duration of each condition, we collected the performance of participants’ classifiers every half-minute in order to understand how well each system enables rapid initialization and iterative improvement over time. After each condition, we collected participants’ self-reported usability ratings of the system in that condition to understand their experiences and challenges that they encountered. For 13 participants, we additionally conducted a semi-structured interview after the experiment was over to understand their challenges in creating and iterating personal classifiers. We further describe our system implementations and experiment design below.
@@ 3.2.1. Recruitment and Participants
We recruited 37 participants with no programming experience by advertising a call for participation on mailing lists of non-Computer Science departments at two major U.S.-based universities. We only selected participants who self-reported having little knowledge of programming and algorithms. There were 27 females, 9 males, and one participant who preferred not to disclose their gender. Most participants were pursuing their bachelor’s degrees except four pursuing more advanced degrees. Regarding political stance, 16 participants identified as liberal, 12 as moderate, and the remaining 9 were evenly distributed among “very liberal,” “conservative,” and “prefer not to disclose” groups. In terms of generative AI usage, 14 participants used it at least every few weeks, 14 participants every few months, and 9 participants rarely.
@@ 3.3.1. Study Design and Procedure
We employed a within-subjects design with the three experiment systems described above. We aimed to obtain at least 0.80 power to detect an effect size of 0.25 in precision at a standard 0.05 alpha error probability using F-tests. The required number of participants for each condition was 28 (Faul et al., 2007). Therefore, our experiment with 37 participants was sufficient for detecting the expected effect. The final experiment protocol was designed iteratively through five pilot experiments to ensure its effectiveness. This study was reviewed by our university IRB and deemed exempt.
@@ 3.3.2. Evaluation Measures
3.4. Data Analysis
3.4.1. Quantitative Modeling
Quantitative results were analyzed using a linear mixed-effects (LME) model, where experiment systems were treated as a fixed effect and participants as a random effect. The dependent variables in our model included classification performance, creation speed, and various subjective measures of user experience. We did not include the order in which participants used the three systems as another fixed effect, as no significant differences were observed for this variable. For each dependent variable, we calculated pairwise differences between the three experiment systems. Finally, we provide basic information about the collected ground truth dataset. Each participant annotated 100 randomly sampled comments from the test dataset, with an average of 65 negative and 35 positive labels per participant. We also conducted a sanity check to confirm that participants exhibited diverse moderation preferences. We found that half of the ground truth comments had, at most, a 75% majority consensus.
3.4.2. Qualitative Coding
4. Results
@@ 4.1. What strategies enabled participants to create high-performing classifiers quickly?
. Figure 6 compares the classification performances across the three systems over 15 minutes, with the endpoint of each line representing the final performance. Table 1 presents the pairwise differences in the final performances among the three systems. We do not observe significant differences in accuracy between the systems. In terms of precision, Rule System showed significantly higher precision than Label System (Est.Diff = 0.097, $p$ 0.05) but did not differ significantly from Prompt System. However, Prompt System showed significantly higher recall compared to both Label System (Est.Diff = 0.205, $p$ 0.001) and the Rule System (Est.Diff = 0.232, $p$ 0.001). Consequently, Prompt System also achieved a significantly higher $F_{1}$ score than the other two systems (Est.Diff compared to Rule System = 0.105, $p$ 0.01; Est.Diff compared to the Label System = 0.141, $p$ 0.001). We focus on $F_{1}$ as a performance metric of interest over accuracy due to the unbalanced nature of participants’ labels.
@@
Table 2 and Table 3 present the pairwise differences in classifier performance at 5 and 10 minutes respectively. Consistent with our earlier findings, the Prompt System demonstrated significantly higher recall and $F_{1}$ scores than the other two systems. Meanwhile, the Prompt System’s recall exhibited a distinct pattern: a rapid increase in the first five minutes, followed by a steady climb in the subsequent five minutes, and a plateau in the remaining time. In contrast, the Rule System showed a consistent increase in recall throughout the entire 15-minute period, whereas the Label System showed a less noticeably upward trend in recall over time. On average, participants reached 95% of their peak recall by 7 minutes, and that of their peak $F_{1}$ score by 5 minutes. We also observed that participants made fewer changes to their prompts as time went on, while they continued actively labeling examples or curating keywords for their rules in the other two systems. These findings suggest that writing prompts can facilitate rapid initialization but may be less effective in supporting iterative improvements.
@@
The precision of the Prompt System and the Rule System remained comparable throughout the experiment. While both the Prompt System and the Rule System showed notable improvements in recall by the end of the experiment, their precision either plateaued or even decreased over time. For instance, the Rule System experienced a gradual decrease in precision during the first 5 minutes, followed by a steady state for the remaining 10 minutes. The Prompt System, despite a slight increase in precision in the initial 2 minutes, subsequently experienced a minor decline before reaching a plateau. These observations align with participant behaviors: participants tended to add more categories of unwanted content rather than increase the specificity of existing rules or prompts during the task because the latter was often considered more cognitively demanding. Although the Label System demonstrated a rapid increase in precision during the first five minutes, it began with such low precision that its final precision was still below that of the other two systems.
@@ 4.2. What strategies did participants find easiest for communicating their preferences?
Unlike professional moderators, social media users often lack the time or opportunity to articulate their moderation preferences clearly (Jhaver et al., 2023). Instead, their moderation decisions are often a result of their intuitive feelings about the content they encounter. Some participants also lacked “a holistic view about what were the alarming comments” (P2), since new content continuously flooded their social media feeds. Labeling examples thus provided them with a natural way to express their intuitive preferences to the algorithm, whereas writing prompts or authoring rules forced them to distill their intuitive preferences into high-level patterns. Participants tended to trust algorithms to infer their preferences from their labels more than they trusted themselves to accurately convey their preferences, especially in time- and attention-constrained settings of personal moderation. As P2 explained, “I would actually trust the algorithm finding trends more [than] me looking at everything and trying to build my own…especially with my lack of experience and lack of time.
@@ 4.3. What strategies were easiest for iterating on classifiers after initialization?
Participants had a clearer understanding of why the Rule System made incorrect predictions than the other two systems. Such transparency helped participants improve their classifiers in some cases, like when they forgot to include spelling variants or similar phrases to those that they explicitly included in their rules. But in other cases, transparency did not lead to iterative improvements because rules could not accommodate contextual variances, as documented in prior research (Song et al., 2023; Reddit, 2023). Our survey and interviews suggest that participants had a slightly better understanding of why their classifiers made mistakes for the Prompt System than for the Label System (Figure 7). Participants found that they could at least review and reflect on problematic prompts in the Prompt System, whereas they could not pinpoint specific labels that led to classification mistakes in the Label System. Moreover, the Label System offered few opportunities for targeted iteration, which were particularly useful as participants found it easier to point out what the classifier should learn from its mistake. This highlights the need for more flexible approaches at different stages of classifier creation, as recent studies advocated (Piorkowski et al., 2023; Daly et al., 2021).
@@
While participants could quickly build classifiers with decent performance using the Prompt System, they found it challenging to incorporate more nuances when iterating on their prompts. Participants rated it as comparably difficult to correct with the Rule System (Figure 7). Many participants described a disconnect in how humans and LLMs perceive the same prompts. They observed that “descriptions could be very subjective” (P4) and asked, “What has [the LLM] previously been trained on? For words they labeled as extreme, I might not think as extreme” (P13). Consequently, many noticed that LLMs seemed to interpret their preferences broadly, causing LLMs to have trouble distinguishing different degrees of a concept. For instance, P13 noted that “[LLMs] have the problem of differentiating between just minor attacks on someone’s character [such as calling someone a fool] and ones that are strong and can be perceived as harm.” Similarly, P8 struggled to teach LLMs to “differentiate between threats and just using the word ‘shoot’ or ‘kill.’” Articulating the fine-grained differences between concepts proved too difficult for many participants.
@@ 4.4. Which strategies did participants prefer overall?
We asked participants to rank the three systems regarding their preferences for using them in real-life content moderation and to explain their rationale for these rankings. Figure 8 illustrates the distribution of participants’ rankings. Surprisingly, despite growing enthusiasm for using LLMs to moderate content, a comparable number of participants still preferred labeling examples or authoring rules to create custom classifiers for content moderation. The Wilcoxon Signed-Rank test showed no significant differences between any of the systems.
@@
  • Whether users want explicit articulation of their preferences. Individual preferences for content curation are nuanced and subject to change. As a result, if users had explicitly stated their preferences in rules or prompts, they would have to continually reflect on their preferences and update their classifiers accordingly. In comparison, the Label System’s algorithms had the potential to learn these subtle changes in preferences from a continuous flux of labels. As P12 explained, “Preferences change over time, and we may not explicitly realize that. So you may accidentally remove some things if your preferences change but you have already explicitly set them in stone. ” However, some participants appreciated the chance to actively reflect on and articulate their preferences through high-level rules or prompts.
@@
  • Whether users can tolerate reviewing many toxic examples. In our study, participants reviewed fewer comments to develop an effective classifier using the Prompt System compared to the other two systems. On average, participants reviewed 133 and 127 examples for the Label System and the Rule System respectively, whereas only 53 examples were needed for the Prompt System. The need to review numerous toxic examples could be distressing, particularly for those creating classifiers to block such content. Whether people are willing to undertake such an emotional toll varies across moderation scenarios. As P8 stated, “If I were moderating for my own sake and not for a community, I would probably just focus more on the ease of creating a filter so that I don’t have to subject myself to these potentially offensive comments.
@@
  • Whether users prioritize precision over recall. Participants demonstrated varied approaches to balancing precision and recall. Some favored precision because they were more concerned about the risk of removing benign comments than approving unwanted content. P12 described their rationale as follows: “ I don’t want to fall into any echo chamber. Even if there are really crazy people saying really crazy things, they make me frustrated, but I still should know about it.” Other participants favored high recall, even if it meant mistakenly removing benign content. P13 expressed this preference: “ Personally, too specific is worse than the too broad. If the filter keeps certain content I really don’t want to see, that’s far worse than if I didn’t get an extra post. Because of how much content there already is, I’ll just get others on my feed anyway.” For these participants, the high recall of the Prompt System is even advantageous during the creation stage. Since LLMs could collect the majority of potentially unwanted examples given initial criteria, participants could focus on reviewing these examples to increase the precision of their prompts.
@@
  • To what extent users value transparency and controllability. An important distinction between the Rule System and the other two systems lies in transparency. As P10 described, “I like rules more because I know what’s going into them. But with the prompts, you don’t know what’s going on. While the prompting is better at catching things, I just like the transparency of rules more.” The Rule System helps users not only understand why exactly a comment is removed; it also enables moderators and users to have clear expectations about which comments will be removed. For users who moderate for a community, this sense of controllability enables them to be more accountable for their moderation decisions.
@@ 5. Discussion
In this work, we compared three prominent strategies for creating custom classifiers, focusing on their support for rapid initialization and easy iteration. Our experiments revealed that, despite its high performance, writing prompts also suffered from challenges in further iteration and lack of transparency. In this section, we discuss how incorporating labeling examples and authoring rules could help mitigate these problems, and how the unique user needs of diverse content curation applications demand different hybrid systems.
@@ 5.1. Hybrid Approaches to Personalized Classifier Creation
Existing tools often fail to accommodate the diversity and fluidity of user needs in content curation. Our experiment indicates the potential to incorporate labeling examples and authoring rules into future content curation systems, thereby providing end users with more flexible strategies to create and iterate on their classifiers.